Find the derivative f m(x) of the following function with respect to x:

fm(x) = ( m

X * n^ x · x ^n )^2

n=1


1
Expert's answer
2020-11-17T13:13:36-0500

For "n=1"

"fm(x)=(mx\cdot1^x\cdot x^1)^2=(mx^2)^2=m^2x^4,"

"fm^{'}(x)=4m^2x^3."